OpenStudy (gabylovesyou):

Let f(x) = 9x - 2 and g(x) = -x + 3. Find f(g(x)).

OpenStudy (chosenmatt):

need help?

OpenStudy (gabylovesyou):

yes please :)

OpenStudy (chosenmatt):

okay sure no problem :)

OpenStudy (chosenmatt):

so did you get anything or come up with anything so far?

OpenStudy (gabylovesyou):

im guessing its f(-x + 3) = 9 * (-x + 3) - 2

OpenStudy (anonymous):

need help guys

OpenStudy (anonymous):

im stuck

OpenStudy (chosenmatt):

i think your right @Gabylovesyou but lemme make sure okay?

OpenStudy (anonymous):

yes that is right clean up with some algebra !

OpenStudy (chosenmatt):

yay!

OpenStudy (gabylovesyou):

what do i do next ?

OpenStudy (chosenmatt):

good job @Gabylovesyou

OpenStudy (anonymous):

\[ 9 (-x + 3) - 2 =-9x+27-2=-9x+25\]

OpenStudy (chosenmatt):

will looks like @satellite73 is helping you on this question but if you need anything just tag me @Gabylovesyou

OpenStudy (gabylovesyou):

@satellite73 how did u get that o-o

OpenStudy (gabylovesyou):

@chosenmatt thank you